# A Mixed-Methods Study on the Implications of Unsafe Rust for Interoperation, Encapsulation, and Tooling

### v3

@@ 6.5. RQ4 - Encapsulation
How do Rust developers reason about encapsulating unsafe code? Interview participants derived requirements for unsafe code from external specifications and the inherent guarantees of the borrow checker. However, they also relied on ad-hoc reasoning and tacit knowledge. Many were uncertain if their encapsulations were sound in all situations. Participants attempted to minimize and isolate unsafe code, and they avoided using unsafe APIs in favor of safer alternatives. However, this was not always possible in certain domain-specific contexts, such as JIT compilers and operating systems, where unsafe code was used pervasively.
Survey respondents who exposed unsafe APIs documented their safety properties, and they relied on documentation to use unsafe APIs correctly. We expect that this documentation could be leveraged by tool designers to create formal specifications for these APIs. When participants chose to expose unsafe APIs, we found that they were motivated by performance, ergonomics, and necessity to the same extent as for any arbitrary use of unsafe code. Respondents rarely refactored unsafe code, even if they used it regularly. Static analysis tools could encourage refactoring by suggesting safe, idiomatic replacements, but it is also possible the the best practices for writing unsafe code make frequently refactoring unnecessary.
@@ 7. Discussion
The majority of our participants attempted to follow several best practices when using unsafe code, such as testing (6.3), writing documentation (6.5), and using safe interfaces whenever possible (6.5). As shown in Table 0, more than 70% of respondents were motivated to use unsafe code when they perceived that there was no safe alternative. However, only 14% (6.4) of these respondents were always certain that it would be impossible to implement a safe alternative. Likewise, 62% (6.5) of respondents who exposed safe APIs for unsafe code were certain “most of the time” that their interfaces were sound, but only 23% were always certain. We cannot rely on developers’ perceptions to accurately reflect the soundness of their code, but this uncertainty suggests that there is still a gap between Rust’s promise of static safety and the reality of how developers use unsafe features.
Multiple factors contribute to this uncertainty. Foreign function calls presented unique challenges due to the differences between Rust and other languages. Rust’s aliasing model is still evolving, so certain features do not yet have an agreed-upon semantics in every possible situation (Jung, 2020). Developers who implement design patterns that sit at the cutting-edge of Rust’s evolving aliasing model will inevitably be uncertain if their implementations are sound. However, all developers who write unsafe code will need a reliable way to determine if their implementations follow the rules of whichever aliasing model becomes canonical. Miri is not entirely adequate for this task, since 61% of the respondents who had used it were deterred by its slow performance and lack of support for key features. However, even a fast, feature-complete version of Miri will not be enough to ensure that developers can always be certain that their implementations are sound, since dynamic analysis is inherently incomplete.
To address the challenges faced by our participants, the next generation of Rust-specific development tools should be capable of providing guarantees of soundness and should support multi-language applications. Static analysis will be particularly helpful for encapsulating foreign functions and guiding developers toward refactoring unsafe code. However, any new approaches to tooling should consider the role of performance and ergonomics as motivations for using unsafe features.
Multi-language Support
Participants who called foreign functions would pass pointers to heap objects in both directions between Rust and other languages. Many used Rust’s memory container types as allocators, and few actively avoided casting foreign pointers into Rust’s safe references. Each of these actions can lead to undefined or undesirable behavior, since foreign code can break Rust’s aliasing rules. However, no production-ready tools are capable of detecting aliasing violations in this context. McCormack et al. (McCormack et al., 2024) were able to extend Miri with support for calling foreign functions, but the limitations of interpretation will prevent their approach from being effective beyond the context of their investigation.
A production-ready tool would need to be capable of instrumenting either native binaries or a shared intermediate format. One promising solution is the Krabcake (Klock and Garza, 2023) project, which aims to extend Valgrind (Nethercote and Seward, 2007) with a plugin for detecting aliasing violations. Valgrind provides a unique advantage compared to other dynamic analysis platforms, since it can instrument native libraries without recompilation. This solution would also be appealing for a significant portion of our respondents, since 44% had used Valgrind. However, Valgrind incurs significant runtime overhead, so it may also be worthwhile to explore developing a similar tool using LLVM’s sanitizer APIs.
Multi-language applications also present a challenge for static verification tools. Several verifiers rely on Rust’s ownership model to provide aliasing information that can simplify the process of specifying higher-order, functional properties (Lehmann et al., 2023; Astrauskas et al., 2019). However, this is only effective when unsafe code is rare. Multi-language applications are exposed to significant sources of unsafety, since C and C++ lack static aliasing restrictions. Our interview participants would benefit from static analysis and inference tools to ensure that static verification remains usable in this context. Techniques such as bi-abduction (Calcagno et al., 2011) would be useful for inferring specifications from unannotated foreign libraries. Contributors to Clang have already proposed a form of lifetime inferences for C++, which would provide annotations that could be used to generate Rust encapsulation with equivalent lifetime semantics (Brænne et al., 2022). Similar analyses would be helpful for determining if foreign types can implement the Send or Sync traits and to warn when functions retain copies of pointers derived from reference-type arguments.
Sound Verification
Despite the need for guarantees of soundness, only 10 survey respondents had used formal methods tools. One possible explanation for the lack of responses to this question is that only Verus (Lattuada et al., 2023) and Kani (VanHattum et al., 2022) had any significant support for unsafe code when we distributed the survey, and Verus had been published only several months prior. Since then, Verus has added support for programs that manipulate raw pointers, and both Gillian-Rust (Élie Ayoun et al., 2024) and RefinedRust (Gäher et al., 2024) have provided alternative methods of verification that specifically target unsafe code. However, none of these tools use formal models that are directly compatible with the aliasing models implemented in Miri, which was used by the majority of our respondents.
To ensure widespread adoption, verification tools could build on the methods and best practices that the majority of our participants already used. Improvements to Miri are likely to be the most effective way to reach our audience. For example, if Miri was extended to support symbolic execution, it provide up to 61% of our survey respondents with guarantees of soundness without requiring them to learn to use a new tool. The Seer (Renshaw and Canciani, 2018) symbolic execution engine was capable of this at one point, but it was created before Miri had support for detecting aliasing violations (Jung et al., 2019), and the project appears to have been abandoned. It would be worthwhile for the Rust community to resurrect this project or implement an equivalent analysis within Miri. Additionally, certain approaches to static verification will inevitably require developers to write specifications, which can be burdensome. However, 90% (6.5) of our respondents who exposed unsafe APIs claim to have already provided informal specifications as comments or documentation. There may be a niche for a lightweight, formal specification language targeted at annotating unsafe APIs with their core safety properties.
Education
Only 6% (6.5) of survey respondents indicated that the Rust community’s guidance and documentation were always enough to determine how to use unsafe code correctly. While the Rustonomicon (Rust Community, 2023b) is an invaluable entry point for learning about Rust’s unsafe features, we expect that its one-size-fits-all approach will inevitably leave certain users wanting more targeted information. We agree with Höltervennhoff et al. (Höltervennhoff et al., 2023) that one of the most impactful ways to improve Rust’s documentation would be to provide a dedicated resource that succinctly summarizes points of contention about Rust’s semantics.
Future interventions should also consider that necessity, performance, and ergonomics can each contribute to a single use case for unsafe code. Understanding these motivations may be helpful for effectively communicating safe alternatives. For example, a profiling tool that integrates information on the distribution and coverage of unsafe code could make it easier to perform a cost-benefit analysis on the tradeoffs between safety and performance. We also expect that lints or static analysis tools could help developers eliminate unnecessary uses of unsafe code, since 88% (6.5) of participants only sometimes refactored their code to remove unsafe features. However, refactoring may not have been necessary for these participants if they had only minimally used unsafe code.

### v4

@@ 5.5. RQ4 - Encapsulation
Interview participants derived requirements for unsafe code from external specifications and the inherent guarantees of the borrow checker. However, they also relied on ad hoc reasoning and tacit knowledge. Many were uncertain if their encapsulations were sound in all situations. Participants attempted to minimize and isolate unsafe code, and they avoided using unsafe APIs in favor of safer alternatives. However, this was not always possible in certain domain-specific contexts, such as JIT compilers and operating systems, where unsafe code was used pervasively. Survey respondents who exposed unsafe APIs documented their safety properties, and they relied on documentation to use unsafe APIs correctly. When our participants chose to expose unsafe APIs, they were motivated by performance, ergonomics, and necessity to the same extent as for any arbitrary use of unsafe code. Respondents rarely refactored unsafe code, even if they used it regularly.
6. Related Work
Here, we discuss relevant results from other studies of unsafe Rust code, comparing our results to quantitative findings where relevant. Most prior studies have examined unsafe code as it is written, either by surveying Rust’s package ecosystem (Astrauskas et al., 2020; Evans et al., 2020; Ozdemir, 2016) and standard library (Cui et al., 2024), or by examining vulnerabilities caused by unsafe code (Qin et al., 2020; Xu et al., 2021; Zheng et al., 2023; McCormack et al., 2025b; Zhang et al., 2023). Other studies have collected blogs, articles, and forum posts to understand the implications of unsafe code for the Rust community. Zeng and Crichton (Zeng and Crichton, 2018) examined a set of articles and comments posted on Hacker News and the Rust subreddit to identify potential challenges to Rust’s adoption. Zhu et al. (Zhu et al., 2022) examined posts on Stack Overflow to determine if developers frequently used unsafe code to resolve compilation errors. These studies provide meaningful contextual information that we can use to cross-validate our findings. However, none can fully answer our research questions.
Far fewer studies have used surveys or interviews to directly engage with developers who use unsafe code. As an extension to their survey of Rust libraries, Evans et al. (Evans et al., 2020) distributed a short survey on the Rust subreddit. They asked developers why they use unsafe code, which operations they use, and what steps they take to ensure that their code is correct; it received 20 responses. Zeng and Crichton (Zeng and Crichton, 2018) also interviewed three Rust contributors to provide additional context in their analysis. Fulton et al. (Fulton et al., 2021) interviewed 16 developers who had advocated for using Rust in industry, followed by a survey that received 178 responses from many of the same communities as ours. Although they used a similar methodology, Fulton et al. (Fulton et al., 2021) mainly focused on evaluating the benefits and drawbacks of Rust as a general-purpose systems programming language instead of examining Rust’s unsafe features. These studies provide evidence that can partially answer a subset of our research questions, vut most do not share our direct focus on unsafe code or do not investigate challenges specific to tool use and interoperation.
Höltervennhoff et al. (Höltervennhoff et al., 2023) are the only other group that we are aware of that has used qualitative methods to study how developers use unsafe code. Their study was published concurrent to our investigation; after we had completed our interviews but before distributing our survey. They interviewed 26 Rust developers who had experience using unsafe code. Except for three who piloted the interview protocol, all of their participants had contributed unsafe code to open source projects included in the unofficial “Awesome Rust,” list (Community, 2026a) of applications. We cover similar topics, but our research questions reflect different, complementary goals. They focused on understanding how developers reason about the security implications of using unsafe code, while we focused specifically on interoperation and tool use. We provide additional details on challenges associated with tool use and interoperation that are missing from their work, and we demonstrate that their results can generalize to a broader population of developers.
Here, we describe how our results compare with these prior studies for each of our research questions.
RQ1 - Interoperation
Interoperation has consistently been one of the most common use cases for unsafe code. In 2020, Astrauskas et al. (Astrauskas et al., 2020) found that 44.6% of all publicly-exported unsafe functions in the Rust ecosystem were static bindings to foreign functions. Other surveys of the Rust ecosystem have had similar findings; Ozdemir (Evans et al., 2020) found that 30% of unsafe blocks were entirely dedicated to foreign function calls, and Evans et al. (Evans et al., 2020) found that 22.5% of unsafe function calls were to foreign functions. Zhang et al. (Zhang et al., 2023) manually analyzed 5,946 unsafe blocks from 140 prominent libraries to determine which unsafe operations were necessary and which could have been avoided. The majority of all raw pointer dereferences and unsafe function calls were marked as necessary to interoperate with foreign libraries. Studies of developers have reached similar conclusions; 70% of Fulton et al. (Fulton et al., 2021)’s survey respondents and all but one of Höltervennhoff et al. (Höltervennhoff et al., 2023)’s interview participants had called foreign functions from Rust. Interoperation was also common among our participants; 70% of survey respondents had called foreign functions.
However, few studies have identified specific challenges associated with interoperation. Cui et al. (Cui et al., 2024) explicitly excluded foreign functions from their investigation of Rust’s safety properties. Xu et al. (Xu et al., 2021) found 12 examples of bugs and vulnerabilities that involved foreign function calls. They were related to application-specific invariants, data races, invalid alignment, and inconsistent layouts. However, the authors provided few details on the nature of these bugs, indicating that they were “straightforward” and not specific to Rust. Fulton et al. (Fulton et al., 2021)’s participants had varying experiences with interoperation, but C++ was particularly difficult due to what one participant described as “rampant aliasing.” Höltervennhoff et al. (Höltervennhoff et al., 2023) indicated that some participants struggled with foreign function calls that involved “memory management,” but that pure functions were easier to reason about. Our interview participants had similar experiences. Calling pure functions was typically straightforward, as long as the bindings were correct and ABI-compliant. When participants struggled to reason about foreign function calls, it was usually because these APIs had aliasing or concurrency patterns that conflicted with Rust’s idioms.
Even fewer studies have identified bugs that meaningfully involved Rust’s aliasing model. In 2023, a set of anonymous authors classified 22 bugs from five popular Rust encapsulations of foreign libraries (Anonymous, 2023). Most of the errors they investigated involved incorrect handling of exceptions and heap objects that were improperly shared across foreign boundaries. They identified two bugs related to Rust’s aliasing restrictions, but they did not describe these bugs in terms of Stacked or Tree Borrows. McCormack et al. (McCormack et al., 2025b) extended Miri to interoperate with an LLVM interpreter and used this hybrid tool to conduct a large-scale evaluation of multilanguage libraries. In Rust, packages and libraries are referred to as “crates” and are published on crates.io. They found 46 bugs from 37 crates, and one of these crates was maintained by the Rust Project. This crate had an aliasing violation that occurred after a foreign API had copied a pointer derived from a “unique” mutable reference. This fits a pattern described by one of our interview participants, who was concerned that foreign APIs would “retain a pointer” from Rust and “keep operating on it later” (P11). Yu et al. (Yu et al., 2025) found additional aliasing violations with CapsLock, an instrumentation tool that uses custom RISC-V hardware capabilities. Their approach performs nearly twice as fast as Miri and has support for inline assembly, but its aliasing model is weaker than both Stacked and Tree Borrows, and the capabilities that it relies on have not been implemented in physical hardware yet. They evaluated their design on a variety of test cases that called foreign functions and found similar categories of aliasing violations as McCormack et al (McCormack et al., 2025b). In particular, certain test cases were “mutating an owned object without borrowing the owner” (Yu et al., 2025).
RQ2 - Tooling
Rust is a relatively new programming language, so most research efforts have focused on developing new tools and not evaluating existing ones. Zeng and Crichton (Zeng and Crichton, 2018) hypothesized that a lack of promotion for tooling could prevent Rust from being widely adopted as a systems language. However, their work was published prior to the creation of most Rust-specific bug-finding tools. Lack of promotion could explain why few of our participants had used formal methods tools. However, several of the tools listed in our survey, such as Verus (Lattuada et al., 2023), had recently been introduced, and few had support for unsafe code. Today, a variety of static verification tools for Rust are being used in industry, and we expect that studying their adoption would be a useful direction for future work.
Höltervennhoff et al. (Höltervennhoff et al., 2023) are the only other group that we are aware of to examine how developers use tooling to assist with writing unsafe code. Certain development tools were somewhat underused among their interview population; only seven of their 26 participants had mentioned using Miri. Since 61% of our survey respondents had used Miri, this does not seem to indicate a broader trend within the Rust community. Likewise, few of their interview participants used automated auditing tools like cargo-audit, while 51% of our survey participants had used one or more of these tools. However, our respondents only needed to have used these tools once, while Höltervennhoff et al. (Höltervennhoff et al., 2023) indicated that participants who audited their dependencies did so “regularly.” Höltervennhoff et al. (Höltervennhoff et al., 2023) also indicated that the majority of their participants had used Clippy, and that some had mixed experiences. Although many of our interview participants had also used Clippy, our investigation was primarily concerned with bug-finding tools, and not linting tools, so we did not highlight it in our analysis or survey. We also cover debuggers, static verifiers, and tools that generate bindings to foreign functions, which were not mentioned in their study.
RQ3 - Motivations
Multiple studies have identified different use cases for unsafe code. Qin et al. (Qin et al., 2020) collected random samples of operations within unsafe contexts from Rust’s standard library and 10 popular third-party libraries. They identified three use cases for these operations; 42% were related to interoperation, 22% were to improve performance, and 14% involved sharing memory between threads. Astrauskas et al. (Astrauskas et al., 2020) used “anecdotal examples” and snippets of source code to create a taxonomy of six purposes for using unsafe code. These included implementing data structures with complex sharing, overcoming the incompleteness of Rust’s type system, emphasizing contracts and invariants, calling foreign functions, using intrinsics or inline assembly, and improving performance. The most frequently used unsafe feature was calling unsafe functions. Fulton et al. (Fulton et al., 2021) found that developers most frequently used unsafe code to call foreign functions, improve performance, and interact directly with hardware or the operating system.
Unlike these studies, we are concerned with “motivations,” which we define as feature-agnostic reasons for using unsafe code. This excludes several of the feature-specific topics mentioned in prior work, such as intrinsics (Astrauskas et al., 2020) or sharing memory between threads (Astrauskas et al., 2020). We cannot reliably assign motivations to specific features without knowing more about the decision-making process behind each use case. Our perspective is similar to Höltervennhoff et al. (Höltervennhoff et al., 2023), who identified necessity, performance, and efficiency as “motivations.” It is unclear whether this distinction was also an explicit consideration for their study. Zhang et al. (Zhang et al., 2023) were able to classify unsafe operations as “necessary” by applying a strict definition: it had to be impossible to implement a safe alternative. In their evaluation of Rust libraries, they identified two reasons why developers chose to use unsafe code when it was not strictly necessary: to improve performance, and because they were unaware of a safer alternative. They did not identify whether developers were aware of a safe alternative but had chosen to use unsafe anyway, since it was easier. Their strict criteria also excludes certain context-specific perspectives on what it means for unsafe code to be necessary. For example, they classified all instances of UnsafeCell as unnecessary, since Rust’s safe RefCell can be used instead. However, it may be necessary to avoid the run-time cost of RefCell in performance-critical applications. Zhang et al. (Zhang et al., 2023) account for this by evaluating the performance impact of several safe alternatives to unsafe operations, but we expect that developers’ subjective experiences will also be helpful in understanding what motivates them to use unsafe code.
RQ4 - Encapsulation
Unsafe code is prevalent throughout the Rust ecosystem, but developers use it minimally and hide it under safe interfaces. In July of 2016, Ozdemir (Ozdemir, 2016) found that 29% of all published crates used unsafe features, but most unsafe blocks were relatively small and tightly-scoped to include only unsafe operations. Evans et al. (Evans et al., 2020) observed similar patterns in September of 2018. The number of published crates had increased by more than ten times, but 29% still directly used unsafe code. Although 38% of crates did not directly use unsafe code, they either directly or indirectly depended on third-party crates that did use these features. In January of 2020, Astrauskas et al (Astrauskas et al., 2020) found that 23.6% of third-party crates used unsafe features. Most instances of unsafe code were relatively straightforward and hidden beneath safe interfaces. The majority of unsafe functions called by libraries were standard (e.g. not closures, function pointers, or trait functions) and defined within the same crate or provided by Rust’s standard library. In particular, 34.7% of all crates had declared unsafe APIs, but they were not visible to users. This aligns with our participants’ experiences; 88% claimed that they would attempt to encapsulate unsafe operations beneath safe APIs, and 61% expected that it would be at least somewhat easy for another developer to understand a random unsafe block or function from their code.
Developers typically mark functions with the unsafe keyword to indicate that users need to uphold a particular safety property. The majority (61%) of the survey respondents who had exposed unsafe APIs only did so when they required an invariant beyond what Rust’s type system could provide. However, Astrauskas et al. (Astrauskas et al., 2020) found that 36.1% of unsafe functions had completely safe implementations. On further investigation, many of these functions had been automatically generated or annotated as unsafe for “legacy reasons.” This could provide an alternative explanation beyond necessity or ease-of-use for why the remaining 39% of our respondents were motivated to expose unsafe APIs. While unsafe functions are relatively common, developers rarely use unsafe traits. Evans et al. (Evans et al., 2020) found that only slightly more than 1% of all crates declared an unsafe trait, and only 6% implemented an unsafe trait. Astrauskas et al. (Astrauskas et al., 2020) had similar findings; only 2.5% of all traits were unsafe, and 40.4% of these declarations originated from five crates. Likewise, Zhang et al. (Zhang et al., 2023) only identified 11 instances of unsafe trait implementations. Astrauskas et al. (Astrauskas et al., 2020) suggested that this lack of use could be caused by a lack of guidance on when a trait should be labeled unsafe. One of our interview participants indicated that certain developers may have misconceptions on the role of unsafe traits. As mentioned earlier in Section 0, one participant who maintained a library that exposed an unsafe trait found that users who had implemented it incorrectly would blame the library, even though “in theory, the burden of proof is on them” (P7). Both the Rust Reference (Rust Community, 2023a) and the Rustonomicon (Rust Community, 2026b) have brief sections on unsafe traits, but additional resources could be helpful.
The safety properties for unsafe features are varied and complex. Cui et al. (Cui et al., 2024) manually reviewed 416 unsafe APIs in Rust’s standard library and identified 19 categories of preconditions and postconditions. Partly for this reason, discrepancies between the signatures of safe APIs and their interior unsafe implementations have been a significant source of bugs and vulnerabilities for the Rust ecosystem. Qin et al. (Qin et al., 2020) examined 170 bugs and vulnerabilities from ten high-profile Rust libraries. Memory safety errors typically occurred in safe contexts but were caused elsewhere by errors in interior unsafe implementations. Qin et al. (Qin et al., 2020) also identified 19 examples of incorrect encapsulations of unsafe code from both the third-party libraries in their sample and within Rust’s standard library. Xu et al. (Xu et al., 2021) exclusively studied registered vulnerabilities and identified similar errors in safe encapsulations. Zheng et al. (Zheng et al., 2023) conducted the broadest empirical study of security issues in the Rust ecosystem to date. They examined 433 unique vulnerabilities that were discovered between 2014 and 2022, identifying 300 affected repositories. Vulnerable components were significantly more likely to include unsafe blocks or functions. The overwhelming majority of all vulnerabilities affected safe functions.
For this reason, it is somewhat concerning that our participants were often uncertain about the correctness of their encapsulations. Höltervennhoff et al. (Höltervennhoff et al., 2023) had similar findings; most of their participants relied on common sense and test coverage to evaluate the security of their code, and few had contributed to projects with formal policies for security or code review. However, many of their participants indicated that they would review unsafe code more carefully to ensure that it upheld all necessary safety invariants. Few participants had ever encountered security issues caused by unsafe code, and none of these issues were critical. Many of our interview participants had encountered memory or thread-safety bugs in unsafe implementations, but few described the security implications of these issues. Security was not a central focus of our study, so it is possible that our participants had reflected on the security implications of their decisions but neglected to mention this during interviews.
@@ 7. Discussion
Our goal in studying interoperation was to describe how developers navigate the differences between Rust and other languages (RQ1) and to determine how tooling could be improved to make this process easier (RQ2). Our interview participants found that interoperation was more or less difficult depending on how the characteristics of foreign APIs matched the inherent expectations of Rust’s aliasing model. When documentation was missing or certain design idioms conflicted with Rust’s aliasing rules, participants had no choice but to minimize their interaction with foreign function calls to reduce the perceived risk of triggering undefined behavior. At the time this survey was conducted, Miri was the only tool capable of finding violations of Rust’s aliasing model—it was both the most popular tool and the most lamented. Participants indicated that they had been deterred from using it due to its lack of support for foreign function calls and inline assembly, or its slow performance.
These challenges are not localized at foreign function boundaries. Developers need to reconcile the semantics of foreign APIs with safe, idiomatic Rust encapsulations, which often requires implementing a layer of interior-unsafe operations. Developers’ motivations for using these operations (RQ3) and their ability to reason about encapsulating unsafe code (RQ4) are inherently relevant to this process. Our population prioritized exposing safe encapsulations whenever possible. Necessity was the most predominant motivation for exposing unsafe APIs, and the overwhelming majority (9) of participants indicated that they would always document any necessary preconditions and postconditions. However, factors that are orthogonal to correctness, like ergonomics and performance, influenced how participants used unsafe code within safe encapsulations. For example, one interview participant indicated that it had been easier to encapsulate a foreign API by transmuting a pointer to a long-lived allocation into a reference with a 'static lifetime. The alternative would have been to derive a context-specific lifetime, and another participant indicated that this pattern could make an API more difficult to use. These motivations were less common but still significant—51% of participants who exposed unsafe APIs cited performance as a motivation.
We expect that Rust developers will continue to prioritize exposing safe APIs, at least in part due to community incentives. More than 80% (9) of participants who had used an unsafe APIs before would prefer a safe alternative at least most of the time, if not always. As one interview participant put it: “everyone wants a safe interface… no one really wants to go use the unsafe heavy ones” (P9). This may explain why interview participants were willing to rely on ad hoc, application-specific reasoning to justify their design decisions in the face of domain-specific challenges and limitations in tooling. As a result, developers are left uncertain about their design decisions. Only 14% (8) of survey respondents who used unsafe code by necessity were always certain that it would be impossible to implement a safe alternative. Likewise, 62% (10) of respondents who exposed safe APIs for unsafe code were certain “most of the time” that their interfaces were sound, but only 23% were always certain.
Participants suggested several potential solutions to this uncertainty. Developers who contributed to applications with a significant body of specialized unsafe features—mainly multilanguage applications, JIT compilers, and operating systems—would benefit from new dynamic analysis tools for finding Rust-specific aliasing bugs at-scale. Static analysis could also be helpful to generate and validate bindings and encapsulations for unsafe functions. However, Rust is also constantly evolving, and interview participants identified several topics that they were uncertain about due to Rust not having a formal, comprehensive specification. For this reason, we expect that our population would benefit from more comprehensive documentation on certain unstable language features.
7.1. Dynamic Tooling
Our participants also indicated that they needed new tools for finding Rust-specific forms of undefined behavior: “a version of Miri that could go across that FFI boundary” (P18), or “…a unique pointer sanitizer” (P1). Miri has experimental support for executing foreign functions from shared libraries. However, it is still several orders of magnitude slower than native execution, prohibiting it from being useful in the types of large-scale C and C++ applications where Rust is increasingly being adopted. We see two potential routes for implementing new tools in this category. The first would be to insert run-time checks for aliasing violations during compilation. This would require lowering a subset of Rust’s type information from the MIR level down into the intermediate representations used by these platforms, so that the instrumentation pass could distinguish between references and raw pointers. Yu et al. (Yu et al., 2025) have had success with this method in CapsLock, but their approach uses an aliasing model that is strictly weaker than Stacked and Tree Borrows. In ongoing work, we are creating BorrowSanitizer (McCormack et al., 2025a), which is implemented without hardware acceleration, allowing us to have parity with Miri’s aliasing models. Our approach uses many of the same APIs as other popular LLVM-based tools.
We expect that compile-time instrumentation tools will provide the combination of performance and compatibility that is necessary to support large-scale, multilanguage applications. However, participants who contributed to JIT compilers would also benefit from dynamic binary instrumentation. Tools in this category, such as Valgrind (Nethercote and Seward, 2007), can encode and instrument native assembly on-the-fly, which would be necessary to support JIT-compiled assembly. In 2023, the Krabcake project (Klock and Garza, 2023) proposed extending Valgrind to be capable of finding Stacked Borrows violations. Several participants had mentioned Krabcake and indicated that it would be a natural extension to their workflows, since they had already used Valgrind with Rust programs. It was also the second most popular dynamic bug-finding tool behind Miri; it was used by 44% of survey respondents. However, Krabcake is no longer actively maintained, and it is not yet capable of finding bugs in real-world programs. The Rust community should consider reactivating this project.
7.2. Static Tooling
Participants also indicated that static analysis tools could be useful for finding undefined behavior across foreign function boundaries. Several analyses could be implemented as extensions to binding generation tools at varying levels of sophistication. One participant wanted a tool that could tell “whether your FFI conforms to your expectations…Do all my symbols match? Do all the types signatures match?” (P5). McCormack et al. (McCormack et al., 2025b) were able to detect errors in foreign function bindings by comparing the size of types in the Rust declarations with the size of types in the LLVM definitions. The authors implemented this as part of a dynamic analysis tool, but their method did not depend on dynamic information, so it could be factored out as a static analysis pass. Changes in Rust’s newest 2024 edition indicate that Rust will continue to evolve toward improving contract visibility at foreign APIs. Rust’s extern blocks now require unsafe annotations (Rust Community, 2024), indicating that developers are responsible for ensuring that foreign function bindings match their definitions. Another participant tried to avoid manually reasoning about the lifetime semantics of foreign APIs “in hopes that in future, there will be a better method for doing this correctly” (P4). In 2022, contributors to Clang proposed adding lifetime annotations to C++, which would be inferred through a whole-program static analysis pass (Brænne et al., 2022). The particular RFC has not had activity since August 2023, but the project was associated with Google’s Crubit analysis tool, which is still in development as of May 2025 (Google, 2025).
Despite their uncertainty, few of our participants had tried using static verification tools. One possible explanation is that when we distributed the survey, only Verus (Lattuada et al., 2023) and Kani (VanHattum et al., 2022) had significant support for verifying unsafe code. Since then, both of these tools have added extensive support for unsafe features. Verus has been used to verify Vest (Cai et al., 2026), a parser combinator library, as well as several other security-critical application components (Verus Team, 2025). Both Kani and Verifast (Jacobs et al., 2025) are being used to verify the entirety of Rust’s standard library as part of a joint initiative between Amazon and the Rust Foundation (Kumar, 2024; Rust Community, 2025g). We expect that these new tools will meet our participants’ needs for stronger guarantees of soundness. For example, two of our interview participants indicated that refinement typing could be useful—“custom subtyping, so that you can specify I want an integer, but only between one and ten” (P4). Flux (Lehmann et al., 2023), a refinement-type checker for Rust, was published concurrent to our study and has since been used to verify several real-world Rust applications (Lehmann et al., 2025; Rindisbacher, 2025). As verification continues to become standardized within the Rust community, it would be worthwhile for future studies to examine how these methods are being deployed in practice.
7.3. Unsafe Code Guidelines
The majority of survey respondents found that the Rust community’s guidance and documentation were helpful, but only 6% (10) indicated that these resources were always enough to determine how to use unsafe code correctly. Interview participants also expressed uncertainty about Rust’s semantics. Some were uncertain about the interaction between Rust’s panic! macro and Drop implementations, with one participant indicating that “real guidelines for library authors” would be helpful in this area. This has been a subject of active discussion since at least 2021 (Jung, 2025), but has yet to be fully documented in community resources (Huss, 2025; Jung, 2025). We expect that other, similar topics lack formal documentation—instead, developers need to examine years-long issue threads within repositories for Rust’s compiler and Unsafe Code Guidelines (Rust Community, 2023b).
We agree with Höltervennhoff et al. (Höltervennhoff et al., 2023) that developers would benefit from having a dedicated resource summarizing points of contention about Rust’s semantics. However, Rust is still a relatively new systems programming language, and certain feature interactions have yet to be fully explored or agreed-upon to the point where writing adequate documentation would even be possible. Developers who implement design patterns that sit at the cutting-edge of Rust’s evolving semantics will inevitably be uncertain if their implementations are sound. Even so, it would be worthwhile for the Rust team to continue to review popular, long-standing issue threads and flag topics that have not been fully documented in the Rustonomicon (Rust Community, 2026b) or the Rust Reference (Rust Community, 2023a). Meanwhile, developers who engage with unsafe code should periodically review relevant issues within the Unsafe Code Guidelines repository to ensure that they are using these features correctly and to have the opportunity to inform compiler developers about their use-cases. However, we still expect that improvements to Miri will have the greatest impact on our survey population, since it was used by a majority of survey respondents, and it serves as somewhat of an executable specification for Rust.